OpenStudy (anonymous):

Maria is walking near a radio transmitter antenna. She begins 3 km west and 4km south of the transmitter, and walks directly toward a point 5 km due north of the transmitter. Maria walks at a constant speed of 4km/hr. During her walk, for how much time was she within 3km of the transmitter. ( Please show work so i cant understand )THanks :)

jimthompson5910 (jim_thompson5910):

were you able to get started at all?

OpenStudy (anonymous):

i got the points down to it

jimthompson5910 (jim_thompson5910):

which points do you have

OpenStudy (anonymous):

(3,-4) (0,5)

jimthompson5910 (jim_thompson5910):

well she actually starts at (-3,-4) because she is west of the transmitter I put the transmitter at (0,0)

OpenStudy (anonymous):

yes yes my bad

jimthompson5910 (jim_thompson5910):

what is the equation of the line through (-3,-4) and (0,5)

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

how about the slope of that line?

OpenStudy (anonymous):

slope is 9/3 =3

jimthompson5910 (jim_thompson5910):

notice how (0,5) is on the line that is the y-intercept

jimthompson5910 (jim_thompson5910):

m = 3 b = 5 so the equation is y = 3x+5

jimthompson5910 (jim_thompson5910):

we'll use that to determine the points of intersection between that line and the circle

jimthompson5910 (jim_thompson5910):

it asks us: " During her walk, for how much time was she within 3km of the transmitter. " so essentially we'll have a circle with radius 3 and center (0,0) and we want to know how long she's been in this 3 km circle

jimthompson5910 (jim_thompson5910):

what is the equation of this circle?

OpenStudy (anonymous):

its x^2+y^2=r^2

OpenStudy (anonymous):

hello?

jimthompson5910 (jim_thompson5910):

sorry about that you are close, but what is r?

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

r is the radius of the circle

jimthompson5910 (jim_thompson5910):

which in this case is 3

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

therefore, x^2+y^2=r^2 x^2+y^2=3^2 x^2+y^2=9 is the equation of the circle

jimthompson5910 (jim_thompson5910):

now let's graph the line y = 3x+5 and the circle x^2+y^2=9 together

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):